### v1

# TrustGeoGen: Scalable and Formal-Verified Data Engine for Trustworthy Multi-modal Geometric Problem Solving

Abstract
Mathematical geometric problem solving (GPS) often requires effective integration of multimodal information and verifiable logical coherence. Despite the fast development of large language models in general problem solving, it remains unresolved regarding with both methodology and benchmarks, especially given the fact that exiting synthetic GPS benchmarks are often not self-verified and contain noise and self-contradicted information due to the illusion of LLMs. In this paper, we propose a scalable data engine called TrustGeoGen for problem generation, with formal verification to provide a principled benchmark, which we believe lays the foundation for the further development of methods for GPS. The engine synthesizes geometric data through four key innovations: 1) multimodal-aligned generation of diagrams, textual descriptions, and stepwise solutions; 2) formal verification ensuring rule-compliant reasoning paths; 3) a bootstrapping mechanism enabling complexity escalation via recursive state generation and 4) our devised GeoExplore series algorithms simultaneously produce multi-solution variants and self-reflective backtracking traces. By formal logical verification, TrustGeoGen produces GeoTrust-200K dataset with guaranteed modality integrity, along with GeoTrust-test testset. Experiments reveal the state-of-the-art models achieve only 49.17% accuracy on GeoTrust-test, demonstrating its evaluation stringency. Crucially, models trained on GeoTrust achieve OOD generalization on GeoQA, significantly reducing logical inconsistencies relative to pseudo-label annotated by OpenAI-o1. Our code is available at https://github.com/Alpha-Innovator/TrustGeoGen.
1 Introduction
Geometric Problem Solving (GPS), as a pivotal branch of mathematical reasoning or more broadly speaking, general reasoning, demands tripartite capabilities [40]: 1) extracting spatial relationships from visual diagrams, 2) performing mathematical reasoning in language modality, and 3) maintaining stepwise logical coherence where intermediate conclusions rigorously justify subsequent deductions. The development of trustworthy AGI systems fundamentally relies on reliably annotated multimodal data that not only bridges visual-textual modalities but also preserves closed-loop reasoning chains. This requirement stems from a critical axiom: a geometrically valid solution demands both correct answers and air-tight logical pathways, as human experts reject proofs with leap-of-faith***Here it means an unjustified or unsubstantiated jump in reasoning, where a conclusion is reached with insufficient logical steps or evidence. arguments even when final results are accurate.
Figure 1: TrustGeoGen overview: (1) Constructor builds geometric premises from a base scene by incrementally adding clauses and enforces constraints via a compiler, (2) Reasoner generates rigorously validated reasoning graph from premises by geometric rules, (3) Sampler selects conclusion nodes from the graph and traces reasoning paths to formalize problems and solutions, (4) Translator converts formal language to natural language for metadata. A bootstrap mechanism further enhances data complexity and problem difficulty.
The recent advancement of Multimodal Large Language Models (MLLMs) [30, 22, 28, 4, 24, 14, 8, 20, 40, 27, 37, 11, 45, 12] has revealed promising capabilities in geometric reasoning, demonstrating competence in solving elementary geometric problems. Parallel developments in external formalized expert systems [7, 18, 44, 21] have extended problem solving capabilities to more complex geometric configurations, with highly specialized models [35, 9, 43] even approaching challenges of the International Mathematical Olympiad (IMO) level. However, these achievements are critically dependent on high-quality geometric data, currently in short supply. The existing datasets are mainly derived from manual annotations of secondary school textbook problems [6, 7, 15, 44, 21], and this process is limited in scale and diversity. Although recent integration of MLLMs (e.g. GPT series [2, 23, 24]) for synthetic question-answer generation [46, 14] has partially ameliorated data scarcity concerns, significant challenges persist in ensuring stepwise solution validitySpecifically, automatically generated procedural labels frequently contain latent inconsistencies, such as misapplied geometric axioms or diagram-coordinate mismatches., which may undermine the logical coherence of reasoning chains during model training.
However, current geometric datasets exhibit four deficiencies: 1) Modality Fragmentation: existing datasets suffer from incomplete modality alignment, lacking comprehensive annotations that synchronize images, textual captions, questions, and chain-of-thought reasoning processes within unified instances. 2) Scalability Issues: Most approaches rely on repurposing human-authored problems from textbooks or examinations rather than generating novel, curriculum-aligned challenges. 3) Cost Constraints: Manual annotation requires costly geometry expertise, while generating solutions through MLLMs consumes excessive computational resources. 4) Trust Deficits: Model-generated solutions lack reliable automated verification mechanisms, introducing risks of undetected logical errors or misaligned reasoning steps.
To this end, we develop TrustGeoGen, a formal language-verified geometric data generation engine that autonomously produces geometric problems, diagrams, and stepwise solutions through four integrated components: 1) a Constructor to generate constraint-satisfying problem premises and diagrams, 2) a Reasoner expanding geometrically valid reasoning graphs with rigorous node verification, 3) a Sampler leveraging the algorithm GeoExplore to extract high-quality reasoning paths, which are later converted into problems by pairing initial premises with derived conclusions, and 4) a Translator rendering formal specifications into natural language. Furthermore, we improve the problem complexity through a bootstrapping mechanism in TrustGeoGen, where formally verified geometric premises in prior stages are recursively fed back to the Constructor to iteratively append new premises. This closed-loop process enables systematic escalation of difficulties while maintaining logical consistency. Additionally, we refine the algorithm GeoExplore to simultaneously explore alternative solution paths and self-correcting reasoning trajectories, generating problems with multiple valid solutions and self-reflective traceback data.
Our proposed TrustGeoGen ensures: 1) Modality Completeness with synchronized dual-form (formal & informal) captions, questions, solutions, and diagrams, 2) Scalability in generating geometry problems by bootstrapping mechanisms without reliance on existing data source and 3) Verifiability through formal validation of all solution steps in its integrated reasoner. Based on TrustGeoGen, we develop a dataset GeoTrust containing 200K samples and a compact testset GeoTrust-test with 240 samples. Experimental results demonstrate state-of-the-art reasoning models DeepSeek-R1 [10] and OpenAI-o1 [25] achieve only 45.00% and 49.17% accuracy respectively on GeoTrust-test. Furthermore, training with our strictly verified reasoning step annotation data proves to be more effective compared to using pseudo-labelsHere, pseudo-label means that the solution to a given question is annotated using SOTA MLLMs (e.g. OpenAI-o1 [25]). with unverifiable reasoning processes. Additionally, our training data from GeoTrust enhances model performance on the out-of-distribution (OOD) testset GeoQA [6], highlighting its broader generalization capability. In a nutshell, our contributions are as follows.
  • We develop TrustGeoGen, a formalized geometric data generation engine, producing scalable and multimodal-integrated data (geometric diagrams, formal & natural language captions, questions and solutions) with trustworthy reasoning guarantees via formal verification.
  • By introducing a bootstrap mechanism, we iteratively generate data with more complex premises and longer reasoning chains, enabling automated complexity scaling.
  • Our GeoExplore series algorithms generates multi-solution problems and self-reflective traceback data, enriching diversity while ensuring trustworthiness.
  • We construct the dataset GeoTrust with 200K samples and GenTrust-test testset, dynamically generated without dependency on prior data sources, ensuring unbiased and extensible evaluating capabilities.
  • Experiments demonstrate that existing MLLMs significantly underperform on high-difficulty geometric reasoning tasks, and our verified data exhibit effectiveness and generalization as training source.
2 Methodology
The data engine TrustGeoGen serves as the core for constructing high-quality geometric reasoning datasets, enabling the generation of complex geometric scenes and reasoning paths to support robust and scalable inference.
2.1 Data Engine
In the process of generating multi-modal geometric reasoning data, it employs a rule-based data construction method augmented by a geometric verifier to synthesize high-quality and scalable datasets. This approach ensures that the data is accurate, controllable, and scalable. As shown in Fig. 1, our framework consists of four components: Constructor, Reasoner, Sampler, and Translator.
Constructor starts the data construction process by building geometric scenes. It incrementally adds new conditions to a basic geometric scene, thus constructing increasingly complex scenes. In TrustGeoGen, a basic geometric scene consists of simple geometric shapes (e.g., triangles, circles, parallelograms) and their associated numerical information (e.g., side lengths, radii, angles). When a base scene is created, it contains an initial set of geometric premises $P_{base}=\{p_{i}^{g},p_{j}^{d}\}$, where $p_{i}^{g}$ represents geometric relationship conditions (e.g., $AB//CD$) and $p_{j}^{d}$ represents numerical conditions (e.g., $\angle BAC=30^{\circ}$). Subsequently, the Constructor incrementally adds new premises to the base scene, increasing its complexity while expanding the premise set $P$. Notably, the Constructor requires that each newly added premise must be connected to the preceding scene to ensure the overall geometric scene remains cohesive and connected. By controlling the number of new premises added, the Constructor generates geometric scenes of varying complexity levels. Each constructed scene is then validated by a geometric compiler [31] to ensure the validity of its geometric relationships. Finally, the process yields a geometric scene with a specified level of complexity and an expanded premise set $P=\{p_{i+m}^{g},p_{j+n}^{d}\}$.
Reasoner leverages a predefined set of geometric theorems to infer new states from premises generated by the constructor, formulating a reasoning graph:
$\small\mathcal{G}=(S,S_{0},R,\hookrightarrow)$ (1)
  • $S$ is the set of states, where each state $s\in S$ represents a derived conclusion or fact in the reasoning process.
  • $S_{0}=P\subseteq S$ is the set of initial states (the premises $P$ generated by the constructor).
  • $R$ is the set of rules, where each rule $r\in R$ defines a logical inference relationship between states.
[…]
Sampler operates on the reasoning graph $\mathcal{G}$ by sampling a target state $s_{n}$ and subsequently searching for its path:
$\small\mathcal{P}=\{(S_{i-1},r_{s},s)\mid\forall s\in S_{i},S_{i-1}\xhookrightarrow{r_{s}}s,i=n,\dots,1\}$ (3)
where each triplet $(S_{i},r_{s},s)$ denotes a specific transition during the reasoning process, derived by applying rule $r_{s}$ to state set $S_{i-1}$ to produce state $s$. Any state $s_{n}$ within the state set $S$ can be treated as a problem to solve in order to derive a reasoning path $\mathcal{P}$. The sampler is initialized with a state $s\in S_{n}$ to retrieve the corresponding transition rule $r_{s}$ and the upstream dependent states $S_{i}$. This process is repeated iteratively until the state set $S_{i}$ consists entirely of initial premises $S_{0}$. At that point, the full reasoning path has been successfully constructed. Notably, during the graph construction process in the Reasoner, once a state $s$ is derived through rule $r_{s}$, no additional state transition relations are appended to $s$. Consequently, the Sampler is always able to identify a unique inference rule $r_{s}$ corresponding to any state $s$, thereby guaranteeing the absence of cycles in the reasoning path.
To ensure the quality of the reasoning paths, the Sampler employs two metrics to filter out unsuitable paths. The first metric evaluates the length of the reasoning path $\mathcal{P}$ against a predefined threshold $\tau_{l}$. If the length of $\mathcal{P}$ exceeds $\tau_{l}$, the path is considered difficult enough. Formally, this condition is expressed as $|\mathcal{P}|\geq\tau_{l}$, where $|\mathcal{P}|$ denotes the number of transitions in the reasoning path $\mathcal{P}$. The second metric assesses the ratio of the states in the reasoning path $\mathcal{P}$ that are derived from the initial premises set $S_{0}$. If this ratio exceeds a predefined threshold $\tau_{r}$, the path is deemed sufficient enough. This condition is formally defined as $\frac{|S_{P}|}{|S_{0}|}\geq\tau_{r}$, where $|S_{P}|$ represents the number of initial premises used in the reasoning path $\mathcal{P}$, and $|S_{0}|$ is the number of all initial premises. By applying these two metrics, the Sampler ensures that only high-quality reasoning paths are retained, thereby maintaining the integrity and usefulness of the generated data. The detailed procedure of the aforementioned reasoning path exploration and filtering algorithm, GeoExplore, is presented in Algorithm 1.
Algorithm 1 GeoExplore
0:  Reasoning graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target state $s_{n}$, threshold $\tau_{l}$, threshold $\tau_{r}$
0:  Filtered reasoning path $\mathcal{P}$
1:  Initialize $\mathcal{P}\leftarrow\emptyset$
2:  $s\leftarrow s_{n}$
3:  while $s\notin S_{0}$ do
4:     Identify rule $r_{s}\in R$ and parent state set $S_{i-1}\subset S$ such that $S_{i-1}\xhookrightarrow{r_{s}}s$
5:     Append transition $(S_{i-1},r_{s},s)$ to $\mathcal{P}$
6:     $s\leftarrow S_{i-1}$
7:  end while
8:  Assess path length: $L\leftarrow|\mathcal{P}|$
9:  Compute premises ratio: $R_{P}\leftarrow|S_{P}|/|S_{0}|$, where $S_{P}\subseteq\mathcal{P}$
10:  if $L\geq\tau_{l}$ and $R_{P}\geq\tau_{r}$ then
11:     Return $\mathcal{P}$
12:  else
13:     Discard $\mathcal{P}$
14:  end if
Translator is responsible for converting the reasoning path $\mathcal{P}$ into a human-readable natural language explanation. To accomplish this, the Translator leverages a large language model (LLM) to ”smooth” the reasoning path. For each state and rule in the reasoning path $\mathcal{P}$, the Translator provides the LLM with a detailed natural language interpretation of the geometric states and inference rules, ensuring that the LLM fully understands the semantic and logical meaning of the components in $\mathcal{P}$. Additionally, the Translator utilizes a few-shot learning strategy to enhance translation accuracy and consistency. Specifically, it includes a pre-curated set of high-quality human-generated translations as few-shot examples, which serve to guide the LLM in generating precise and context-appropriate explanations. By iteratively processing each triplet $(S_{i-1},r_{s},s)\in\mathcal{P}$, the Translator generates a coherent chain-of-thought explanation that mirrors the logical flow of the reasoning graph $\mathcal{G}$. This ensures that the resulting translation not only reflects the exact logical derivations but also facilitates human understanding by articulating complex geometric relationships in an accessible and comprehensive manner.
Algorithm 2 GeoExplore-M
0:  Reasoning graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target state $s_{n}$, threshold $\tau_{l}$, threshold $\tau_{r}$, number of searches $n$
0:  Set of filtered reasoning paths $\mathcal{P}_{\text{set}}$
1:  Initialize $\mathcal{P}_{\text{set}}\leftarrow\emptyset$
[…]
3:  while True do
4:     Initialize $\mathcal{P}\leftarrow\emptyset$
5:     $s\leftarrow s_{n}$
6:     while $s\notin S_{0}$ do
7:        Identify set of rules $\mathcal{R}_{s}\subseteq R$ and corresponding parent state sets $\mathcal{S}^{m}_{i-1}\subseteq S$ such that $\forall r_{s}\in\mathcal{R}_{s},\exists S_{i-1}\in\mathcal{S}^{m}_{i-1}:S_{i-1}\xhookrightarrow{r_{s}}s$
8:        Select $r_{s}\in\mathcal{R}_{s}$ and $S_{i-1}\in\mathcal{S}^{m}_{i-1}$ (choose different options in different searches)
[…]
10:        if $(S_{i-1},r_{s},s)\notin\mathcal{P}$ then
11:           $\mathcal{P}\leftarrow\mathcal{P}\cup(S_{i-1},r_{s},s)$
12:        end if
13:        $s\leftarrow S_{i-1}$
14:     end while
15:     Assess path length: $L\leftarrow|\mathcal{P}|$
16:     Compute premises ratio: $R_{P}\leftarrow|S_{P}|/|S_{0}|$, where $S_{P}\subseteq\mathcal{P}$
17:     if $L\geq\tau_{l}$ and $R_{P}\geq\tau_{r}$ then
18:        $\mathcal{P}_{\text{set}}\leftarrow\mathcal{P}_{\text{set}}\cup\{\mathcal{P}\}$
19:     end if
20:     if used_options contains all possible options then
21:        Break
22:     end if
23:  end while
24:  return  $\mathcal{P}_{\text{set}}$
2.2 Bootstrap Augmentation
As previously described, TrustGeoGen begins by constructing a complex geometric figure from a completely random initial scenario and then derives reasoning paths for this instance. While this approach is scalable, its efficiency in generating high-quality data is relatively low, as it may randomly produce geometric scenarios that fail to yield meaningful or valuable results. To address this, TrustGeoGen employs an iterative bootstrap strategy to guide the generation process, thereby improving the efficiency of producing high-quality data. In each iteration, TrustGeoGen utilizes the metrics from the Sampler to evaluate the quality of sampled data on a given graph $G$. If the average quality of the sampled data on a particular graph $G$ is sufficiently high, TrustGeoGen selects this data as the basis for the subsequent iteration’s initial scenario. Specifically, the process can be represented as $P^{\prime}_{base}=P=\{p_{i+m}^{g},p_{j+n}^{d}\}$, where $P$ serves as the starting point. The Constructor then incrementally adds new premises to this foundation, increasing the complexity of the scenario. This leads to the generation of $P^{\prime}=\{p_{i+m+x}^{g},p_{j+n+y}^{d}\}$. Following this, the Reasoner, Sampler, and Translator collaboratively work to complete graph construction, problem-solving, and translation, ultimately producing a new batch of high-quality data.
2.3 Multi-solution Data
In the field of GPS, exploring multiple solutions to a given problem is a crucial method for understanding geometric relationships in depth. By investigating different solution logics, a model can examine geometric scenarios from various perspectives, achieving a comprehensive understanding of the relationships among geometric elements. TrustGeoGen facilitates the construction of multi-solution datasets for specific geometric scenarios through the graph-building procedure of the modify Reasoner and the solving procedure of the Sampler.
[…]
Algorithm 3 GeoExplore-T
0:  Directed graph $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, target state $s_{t}\in S$, threshold $\tau_{p}$
0:  Trace-back reasoning data $\mathcal{D}$
1:  Initialize $\mathcal{D}\leftarrow\emptyset$
2:  Compute all reasoning paths $\mathcal{P}^{t}_{\text{set}}=\{\mathcal{P}_{1}^{t},\mathcal{P}_{2}^{t},\dots\}$ to $s_{t}$ using Algorithm 2
3:  Randomly sample a state $s_{e}\in S$ such that $s_{e}\notin\text{UpstreamDependencies}(\mathcal{P}^{t})$ for any $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$
4:  Search for a reasoning path $\mathcal{P}^{e}$ leading to $s_{e}$
5:  if $\exists\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}},\text{Overlap}(\mathcal{P}^{e},\mathcal{P}^{t})\geq\tau_{p}$ then
6:     Add $\mathcal{P}^{e}\cup\mathcal{P}^{t}$ to $\mathcal{D}$
7:  end if
8:  return  $\mathcal{D}$
2.4 Self-reflective Traceback Data
Recent studies on the reasoning capabilities of large models suggest that these models acquire substantial knowledge during the pre-training phase. By providing training data with various cognitive templates during the post-training phase, the models can better utilize their inherent knowledge, thereby enhancing reasoning ability and improving response quality[42, 41]. Among these cognitive templates, self-reflection trace-back reasoning serve as highly effective approaches. However, self-reflection and retrospective reasoning often occur during the author’s cognitive process and are rarely preserved in the final outputs, making the collection of relevant data particularly challenging. These types of data are even more scarce in the domain of multi-modal geometric problem-solving.
TrustGeoGen defines trace-back reasoning data as a sub-graph of $\mathcal{G}=(S,S_{0},R,\hookrightarrow)$, characterized as the union of sub-graphs with identical upstream dependencies but differing final states. Specifically, for a given piece of trace-back reasoning data, the process involves first deriving an incorrect final state, then retrospectively navigating a correct reasoning path (shared by both the incorrect and correct paths), and ultimately arriving at the correct final state. To achieve this, TrustGeoGen selects a target state $s_{t}$ and employs Algorithm 2 to enumerate all possible reasoning paths $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$. Next, a state $s_{e}$ is randomly sampled from the state set $S$, ensuring that $s_{e}$ does not belong to the upstream dependencies of any $\mathcal{P}^{t}$. Subsequently, TrustGeoGen searches for a reasoning path $\mathcal{P}^{e}$ leading to $s_{e}$. If $\mathcal{P}^{e}$ overlaps with any $\mathcal{P}^{t}\in\mathcal{P}^{t}_{\text{set}}$ and the overlapping portions exceed the threshold $\tau_{p}$, TrustGeoGen outputs $\mathcal{P}^{e}\cup\mathcal{P}^{t}$ as trace-back reasoning data. Notably, Algorithm 2 is utilized to identify all possible reasoning paths for $s_{t}$ to ensure that $s_{e}$ does not appear in any potential upstream dependencies, thereby maintaining the integrity of the reasoning process leading to $s_{t}$.The details of the trace-back data acquisition algorithm, GeoExplore-T, are presented in Algorithm 3.
Figure 2: Simple examples of Multi-solution and Self-reflective traceback data.
Figure 3: Data distribution and augmentation. (a) Distribution of reasoning lengths, with most samples below 60 steps and a sharp decline beyond 80; the zoomed-in view highlights the ultra-long reasoning range. (b) Distribution of premise ratios, reflecting variability in logical dependencies. (c) Comparison of reasoning lengths before and after bootstrap augmentation, showing increased deep-reasoning samples (reasoning length $\geq$ 40) and reduced shallow ones.
Figure 4: Visualization examples of different difficulty levels in GeoTrust-test, where ”reasoning length” indicates step length of reasoning path and ”premise ratio” refer to the ratio of premises unutilized during formal reasoning and premises provided in the question.
3 Data Analysis
The modality-completed trustworthy geometric reasoning data holds great potential for improving the geometric problem-solving capabilities of MLLMs. Moreover, it could generalize these abilities to other tasks requiring deep reasoning. To validate it, we utilized TrustGeoGen to produce the dataset GeoTrust of 200K samples, from which 8K were sampled as the training set. Additionally, we employed different thresholds, $\tau_{l}$ and $\tau_{r}$, in conjunction with manual screening to curate a testset, GeoTrust-test, with 240 samples of varying levels of difficulty.
3.1 Data Distribution
TrustGeoGen was executed on a 60-core Intel Xeon (32) CPU @ 2.900GHz for two days, generating an initial dataset of 200K geometric reasoning instances in formal language. During the data generation process, we set the two filtering thresholds in the Sampler to $\tau_{l}=5$ and $\tau_{r}=0.5$, respectively. The distributions of the reasoning length and premises ratio for the resulting 200K dataset are illustrated in Fig. 3(a) and Fig. 3(b), respectively. The majority of the data is concentrated in the region where the reasoning length is less than 60, and it decreases sharply beyond 80. However, a small portion of the data exhibits reasoning lengths exceeding 150, indicating a considerably high level of reasoning complexity. For a more detailed view of the distribution, please refer to the zoomed-in section of Fig. 3(a).
Table 1: Performance of state-of-the-art multi-modal language models on GeoTrust-test, which is divided into four difficulty levels.
Model #Params Total (out of 240) $Tier_{1}$ (out of 60) $Tier_{2}$ (out of 60) $Tier_{3}$ (out of 60) $Tier_{4}$ (out of 60)
num acc num acc num acc num acc num acc
LLaVA-1.5 [19] 7B 7 2.91% 3 5.00% 2 3.33% 0 0.00% 2 3.33%
Qwen2-VL-Instruct [36] 7B 9 4.17% 5 8.33% 2 3.33% 0 0.00% 2 3.33%
Qwen2.5-VL-Instruct [5] 7B 24 10.00% 12 20.00% 6 10.00% 3 5.00% 3 5.00%
GPT-4o [24] - 55 22.92% 25 41.67% 15 25.00% 9 15.00% 6 10.00%
Claude-3-7-Sonnet [3] - 83 34.58% 35 58.33% 17 28.33% 24 40.00% 7 11.67%
DeepSeek-R1 [10] 671B 108 45.00% 40 66.67% 24 40.00% 25 41.67% 19 31.67%
OpenAI-o1 [25] - 118 49.17% 42 70.00% 28 46.67% 26 43.33% 22 36.67%
3.2 Deep Reasoning Augmentation
As mentioned earlier, the randomly constructed scenarios by TrustGeoGen exhibit a rapid decline in reasoning length within the ultra-long range, posing challenges for generating deep-reasoning problems. To address this, we selected 226 samples with the longest reasoning lengths from an initial dataset of 200K and applied bootstrap augmentation to increase the proportion of deep-reasoning problems. Finally we obtained 376 geometric scenes after applying the bootstrap method. As shown in Fig. 3(c), after the bootstrap process, the data distribution in the region where reasoning length is smaller than 40 decreased, while it increased in the region where reasoning length is larger than 40. This method can be repeatedly applied to efficiently construct reasoning problems with significant depth and complexity.
3.3 Construction and Analysis of GeoTrust-Test
The synthetic data generated by TrustGeoGen can also be utilized to evaluate a model’s capabilities in the domain of deep geometric reasoning. From an initial dataset of 200K samples, we manually curated a set of 240 high-quality problems as the test set, graded by different levels of difficulty. As shown in Fig. 4, these problems are divided into four tiers, with each tier containing 60 problems. The reasoning lengths for $Tier_{1}$ range from 5 to 10 steps, $Tier_{2}$ spans from 10 to 20 steps, $Tier_{3}$ ranges from 20 to 50 steps, and $Tier_{4}$ exceeds 50 steps. It is worth noting that, to introduce distractors into the questions, the premises provided may not necessarily all be utilized in the reasoning process, resulting in a potentially lower premise ratio.
4 Experiment
To validate the quality and effectiveness of the geometric Problem Solving data generated by TrustGeoGen, we conducted experiments to explore:
  • Whether existing MLLMs demonstrate competent performance on complex geometric problems?
  • Whether geometric reasoning data with rigorous formal verification improve MLLMs’ capabilities and provide specific advantages?
  • Whether GeoTrust, constructed without prior data sources, can generalize to OOD geometric testset?
4.1 Dataset, Metric, and Implementation Detail
Dataset. To evaluate the capability of MLLMs in solving complex geometry problems, we construct GeoTrust-test by partitioning the original GeoTrust dataset. As introduced in Sec. 3.3, the testset comprehensively covers four difficulty levels ranging from $Tier_{1}$ to $Tier_{4}$. To validate the effectiveness of our data construction, we additionally sample 3K geometry problems as GeoTrust-train, with each difficulty tier ($Tier_{1}$ to $Tier_{3}$§§§Considering the maximum output token limit of the trained open-source models, we excluded data from $Tier_{4}$.) containing 1K instances. GeoQA [6] is employed to evaluate the OOD generalization performance, comprising 4,998 geometric problems and solutions in natural language.
Metric. To eliminate potential evaluation bias introduced by the multiple-choice format, all input questions in the experiments are presented without answer options. Both the model’s final output answers and GT labels are extracted and converted into floating-point values for numerical comparison, with a relative error tolerance of 1% permitted.
Implementation Detail. In the experimental training protocol, all models are trained through supervised fine-tuning (SFT) with full-parameter updates for one epoch to prevent overfitting. All training and evaluation processes are conducted on 8 NVIDIA A100 (80G) GPUs.
4.2 Performance of Existing MLLMs
As shown in Tab. 1, the experimental results from the GeoTrust-test testset reveal critical insights into the capabilities of MLLMs in tackling complex geometric problems. Existing open-source models, such as Qwen2.5-VL-7B-Instruct [5], demonstrate significant limitations, achieving only a 12.92% overall accuracy (31/240), which also underscores the inherent challenges of the dataset and confirms its independence from prior biases in existing open-source testsets. Besides, closed-source commercial models like OpenAI-o1 [25] exhibit stronger reasoning capabilities, solving 109 problems (45.42% accuracy), though their performance sharply declines on higher-difficulty $Tier_{4}$ questions (18.33% accuracy), highlighting persistent gaps in handling advanced geometric reasoning. Notably, the open-source reasoning model DeepSeek-R1For DeepSeek-R1, we only provide natural language questions. [10] matches OpenAI-o1’s overall accuracy, bridging the performance gap between open and closed-source systems and emphasizing the potential of scalable architectures. A consistent trend across all models is the progressive degradation of accuracy as problem difficulty escalates, validating the rigorous tiered difficulty framework of GeoTrust-test. These findings collectively emphasize the need for enhanced reasoning architectures and training paradigms to address the steep complexity gradient in geometric problem solving.
4.3 Advantages Provided by Trustworthy Data
Table 2: Comparison of performance on GeoTrust-test using trustworthy GeoTrust-train data and pseudo-labels annotated by OpenAI-o1 [25]. All trained on Qwen2-VL-7B-Instruct [5].
Training Data Total $Tier_{1}$ $Tier_{2}$ $Tier_{3}$
- 7 5 2 0
pseudo-labels 21 (+14) 13 (+8) 3 (+1) 5 (+5)
100% GeoTrust-train 38 (+31) 27 (+22) 6 (+4) 5 (+5)
50% GeoTrust-train 27 (+20) 22 (+17) 4 (+2) 1 (+1)
30% GeoTrust-train 12 (+5) 10 (+5) 2 (-) 0 (-)
To validate the effectiveness of the trustworthy data generated by TrustGeoGen, we compared our data with the ones generated by OpenAI-o1 [25]. We collect pesudo-labels generated by the OpenAI-o1 [25] via prompting the model with the problems in GeoTrust-train. We train Qwen2-VL-7B-Instruct [36] on both GeoTrust-train (with 100%, 50% and 30% data) and pesudo-labels, and report the number of problems solved from $Tier_{1}$ to $Tier_{3}$. As shown in  Tab. 2, the model trained on 100% of GeoTrust-train outperforms the model trained on pesudo-labels, which demonstrates that trustworthy data hold substantial advantages over SOTA-generated pseudo-labels under equivalent data amount. Besides, as the amount of GeoTrust-train data increases, the model’s overall performance continues to improve, which further validates the effectiveness of our dataset. However, while the improvements on $Tier_{1}$ are notable, the gains on $Tier_{2}$ and $Tier_{3}$ remain significantly less pronounced, suggesting that supervised fine-tuning (SFT) does not produce substantial effects on more challenging geometric reasoning tasks.
4.4 OOD Generalization
The experiments are conducted on three pre-trained vision-language models: LLaVA-1.5-7B [19], LLaVA-1.5-13B [19] and Qwen2-VL-7B-Instruct [36]. For each model, we evaluate on GeoQA [6] with two training settings: (1) fine-tuning solely on the GeoQA dataset, and (2) fine-tuning with a combination of GeoQA and GeoTrust-train datasets to explore generalization. The results in Tab. 3 demonstrate that incorporating GeoTrust-train data not only achieves substantial improvements over the original pre-trained models (e.g., +11.81% for LLaVA-1.5-7B and +5.57% for Qwen2-VL-7B ), but also surpasses models fine-tuned solely on the in-distribution GeoQA training set. For example, LLaVA-1.5-7B improves from 25.73% accuracy (GeoQA-only fine-tuning) to 28.12% when augmented with GeoTrust-train, marking a +2.39% absolute gain over the in-distribution baseline, highlighting the generalization capability of our data.
Table 3: OOD Generalization on GeoQA [6] when GeoTrust-train is leveraged for training, where $\Delta$ indicates the improvement compared to the pre-trained baseline.
Model Training Data acc $\Delta$
LLaVA-1.5-7B [19] - 16.31% -
GeoQA 25.73% +9.42%
GeoQA+GeoTrust-train 28.12% +11.81%
LLaVA-1.5-13B [19] - 18.83% -
GeoQA 29.84% +11.01%
GeoQA+GeoTrust-train 33.16% +14.33%
Qwen2-VL-7B-Instruct [36] - 37.67% -
GeoQA 41.11% +3.44%
GeoQA+GeoTrust-train 43.24% +5.57%
5 Conclusion and Further Discussion
Conclusion. We present TrustGeoGen, a multimodal-integrated and scalable geometric data engine that automatically generates fully verified geometric reasoning data. The resulting GeoTrust dataset demonstrates validated effectiveness and generalization capabilities, with its trustworthy reasoning data showing significant advantages over conventional pseudo-labels. Notably, the GeoTrust-test exposes critical limitations of existing MLLMs in handling complex geometric reasoning tasks. GenGen represents a crucial first step towards trustworthy geometric problem solving, establishing a new paradigm for generating rigorous data.
Trustworthy Geometric Problem Solving. Models should produce verifiable reasoning steps, not merely correct answers. Achieving this demands both reliable data and robust evaluation mechanisms are required. Leveraging the generated data from TrustGeoGen, further study can focus on autoformalization approachs that translate natural language reasoning into formalized steps, enabling automated verification of each deduction’s logical validity. Moreover, TrustGeoGen can dynamically generate evaluation data to prevent prior data leakage during evaluation, ensuring higher reliability in unseen geometric scenarios.
Formal Enhanced Reasoning. TrustGeoGen’s formalized reasoning environment enables the generation of trustworthy geometric data by constructing rigorous reasoning graphs that ensure logical correctness at each step. These graphs provide a structured foundation for exploring diverse mathematical reasoning strategies through tailored sampling methods: the current work implements multi-solution and self-reflection traceback data, while future extensions could incorporate the idea of reverse thinking and categorical discussion, etc. These ideas could be further enhanced by integrating alternative training approaches (e.g. RL) to generalize to other reasoning scenarios.
Appendix
Appendix A Related Work
Advances in Multimodal Large Language Model.
Recent advances in Large Language Models (LLMs) have demonstrated remarkable progress in linguistic intelligence, achieving human-level capabilities in various applications [26, 33, 34, 32]. Building upon this foundation, the research community has focused on extending these text-based architectures to process visual information, leading to the emergence of sophisticated Multimodal Large Language Models (MLLMs) [4, 1, 29]. These integrated systems typically employ modality alignment mechanisms, such as Q-former [17] and linear layers [19], to establish connections between visual representations and text embeddings. Although current MLLMs show strong performance in standard vision-language benchmarks [13, 39, 38, 22, 16], their effectiveness decreases significantly when processing mathematical visualizations that require reasoning. Recent specialized approaches address this limitation through targeted training strategies. For example, geometric reasoning capabilities have been enhanced through domain-specific datasets containing annotated diagrams [14, 46].
Model Name Model / API Version
GPT-4o [24] gpt-4o-2024-11-20
Claude-3-7-Sonnet [3] claude-3-7-sonnet-20250219
DeepSeek-R1 [10] deepseek-r1
OpenAI-o1 [25] o1-2024-12-17
Table 4: Model / API versions used for evaluation across different MLLMs.
Geometric Problem Solving.
As a challenging task, geometric problem solving requires understanding diagrams, interpreting symbols, and performing complex reasoning. While MLLMs have shown remarkable proficiency in human-level reasoning and generation capabilities, they struggle with automatic geometric problem solving due to their pre-training on natural images and texts, as well as the lack of automated verification in the problem-solving process. To address this limitation, researchers have proposed various approaches to improve the understanding of geometric images and corpora, including unimodal pre-training, vision-language alignment, and visual instruction tuning [40, 46, 14, 6, 7, 44, 15, 16]. Another major approach involves the use of formalized solvers to tackle geometric problems. While these solvers [35, 9, 31] are capable of addressing challenges at the level of the International Mathematical Olympiad (IMO), they require precise modeling of geometric problems to fit the solver’s language. This necessity introduces significant obstacles in terms of practical application and generalization, limiting the usability in real-world scenarios.
Appendix B Versions of Closed-source Model
During the evaluation on GeoTrust-test, we utilize some closed-source commercial models, with the specific version detailed in Table 4.
Appendix C More Visualization Example
In Fig. 5, we provide one more visualization example in GeoTrust with 94 reasoning steps and 0.95 premises ratio.
Figure 5: One visualization example in GeoTrust.